Dancingwithdeath:

Where is the removable discontinuity of f (x) = StartFraction x + 5 Over x squared + 3 x minus 10 EndFraction located? x = –5 x = –2 x = 2 x = 5 This is a pre-calculus question for grade 12 I just need some help!

Luigi0210:

A removal discontinuity can be defined as a point on a graph that is undefined. They can exist if there is a blip in the function Or if there is a common factor within the denominator/numerator, in your case: \[\Large f(x) = \frac{ x+5}{x^2+3x-10}\] can be factored into: \[\Large f(x) =\frac{x+5}{(x+5)(x-2)}\] As you can see, there is a common factor of x+5 You just gotta set this equal to 0 to find the point of discontinuity: \[\large x+5 =0 \]